From GlueXWiki
Jump to: navigation, search

August 7, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status (Mike, Beni)
  3. FDC status interaction event reconstructed with cluster timing algorithm(Lubomir)
  4. Electronics (Chris, Nick)
  5. Engineering (Bill)
    • Chiller
    • Cathode scan [1] (Luke)
  6. Commissioning Wiki Page (all)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Sean (NU), Bill, Luke, Mike, Chris, Nick, Simon, Beni, Eugene, and Lubomir (JLab).


- Beni: the synchronized run (CDC +BCAL) didn't work over the weekend. Something is wrong with the data, as if the the signals were not in the time window. Still, the DAQ was running over the weekend without crushing. Will try it again this weekend.

CDC update

- Beni: the noise was studied with modified HVB in which the resistors, that reduce the signal by 1/2, were removed. As result, both the noise and the signals increased by factor of 2. Beni will try also with the FDC pre-amp.

- Beni and Mike tried to improve the ground using: grounding braid connected to one HVB, Al foil covering the whole magnet opening. No improvements so far, but maybe one has to ground many more HVBs before you see any effect (Beni). More discussions about the noise: what is the resistivity of the CDC aluminum ring and was it anodized, what kind of grounding straps can be attached and how.

FDC update

- Lubomir was improving the tracking: significant advantage if one uses the timing of each ionization cluster (the positions of the maximums in the fADC data). The time of each cluster should be the same for the upper, down strips and the wires. This allows to reconstruct multiple hits (ionization clusters) on the same wire or strip and separate tracks very close to each other, as demonstrated on the attached plot (shown are avalanche positions, no drift time correction yet).


- Casey installed almost all CDC retainers (the retainer keeps the connector to the pre-amp in place).


- Bill: the compressor will come today and will be replaced tomorrow. The chiller failed last Saturday due to compressor overheating, on Monday technicians from facilities found that the compressor has to be replaced. Bill: the most important question is why it failed and do we need spare chiller or parts for the long run. Eugene: during runs we don't want to wait a week for the chiller to be repaired. Let's first see if this failure was an exception, but in any case having a spare compressor (~$1k) will not be a problem.

- Bill: the water flow/quality is a suspect; rust found in the water filter. Therefore a flow meter for the chilled water was installed and Dave should be able to read it. According to Tom, we are going to change the water filters every month. Eugene: we may need better (centralized) filtering system. Bill asked Carroll Jones if the water quality can be improved.

- Luke did already measurements in three positions and is working now on the fourth. The averaged (over 8 strips) pitch values as function of their positions are shown on the 2D plot. They vary between 5.009 and 5.014 mm (color coded). Eugene: to describe this with a model there are central regions for each foil you can consider flat, and dips in between where the Kapton tape is glued.


- The estimates are that at 10^-4 R.L. radiator and 50nA electron beam we will have 180Hz charged particle from 2mm plastic target. The fraction that goes on the FDC has to be simulated, but if we assume 20Hz charged particles on the FDC, we will need about 2 beam-on-target days to have ~3M events. Based on the experience with the package 3 this will be enough for the alignment in one configuration, so in total 4 days (magnet on/off). The target has to be placed as far as possible which is at ~4m upstream from the face of the CDC, at the exit of the vacuum pipe there.

- Eugene: be aware that for the commission we will have ~3m air in front of the solid target (in nominal position) which is ~0.8% R.L. while the 2mm target is ~0.25%. We may decide use carbon or Al target instead of plastic. Simon: for the CDC we will use the target at nominal position. It should not be a problem to have several targets at the same time.

- It is critical to have the sparsification firmware ready by the middle of September so that Beni has time to implement it by the beginning of October for the commissioning run. If the whole firmware is not ready we need at least the "pulse mode" where all the samples are read out, but only for channels with signals above a threshold. In this case we need to think about efficient trigger, like FCAL and TOF coincidence.